OpenStudy (anonymous):

A ferris wheel with a diameter of 110 feet makes a 1/2 of a turn. You are sitting in a seat on the outer rim of the ferris wheel. How far did you just travel?

OpenStudy (callisto):

=1/2*(2πr) =π(110/2) = 55 π not sure for that